can ypu please solve
tan(2tan-1 1/5)
tan-1(1) +cos-1(-1/2)
 
 

Grade:12

2 Answers

Utsav Basu
6 years ago
Ans 1: 
Let θ = tan-1 1/5
Then the expression reduces to
   tan 2θ
= 2tan θ/(1 – tan2θ)
= 2tan(tan-1​ 1/5)/(1 – (tan(tan-1​ 1/5))2)
= 5/12
 
Ans 2:
   tan-1(1) +cos-1(-1/2)
= π/4 + 3π/4
= π
Rohit Sharma
6 years ago
1- Let y= tan-1(1/5)
     Hence, tan y=1/5
     so the question would be tan2y which can be expanded and the value of tany should be substtuted and the final answer               would   be 5/12
2- tan-1(1) + cos-1(1/2) = Pie/4 + 3pie/4 = Pie